OpenStudy (anonymous):

HELP PLEASE! Nelson wants to use a few pieces of silver to make a bracelet. Some of the jewellery is 70% silver, and the rest is 58% silver. Nelson needs 26g of 65% silver for the bracelet. How much of each alloy should he use?

OpenStudy (anonymous):

let x=grams of 70% silver let y= grams of 58% silver x+y=26 0.70x+0.58y=0.65 i just need to know, is that the right way to set it up?

OpenStudy (anonymous):

that's almost right. your last equation should read 0.70x+0.58y=0.65(x+y)

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

uhmm.. oh yess i think i get why you put the (x+y) at the end

OpenStudy (anonymous):

yeah cause you need 26g of the 0.65

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

couldnt you just write 0.70x+0.58y=0.65(26)?

OpenStudy (anonymous):

yeah that works as well :)

OpenStudy (anonymous):

k thanks so much youre saving my life here :)

OpenStudy (anonymous):

i'm glad to be of service :)

OpenStudy (anonymous):

This one !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got x=-35.5

OpenStudy (anonymous):

I don't think that is right

OpenStudy (anonymous):

because it shouldn't be a negative :P

OpenStudy (anonymous):

no you should get 15.167

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

Here ill show you my steps

OpenStudy (anonymous):

Awesome

OpenStudy (anonymous):